If oue population of men, height is normaly distributed with 70 inches(5"10") being the mean height. What is the probability that a randomly selected man will be 74 inches(6"2)or taller with the standard deviation is 2 inches.
Answer by stanbon(75887)

If oue population of men, height is normaly distributed with 70 inches(5"10") being the mean height. What is the probability that a randomly selected man will be 74 inches(6"2)or taller with the standard deviation is 2 inches.
---
This problem was posted and answered yesterday.
----
z(74) = (74-70)/2 = 2
P(x > 74) = P(z > 2) = normalcdf(2,100) = 0.0228
===================
Cheers,
Stan H.
============